OpenStudy (anonymous):

Convert these unlike fractions to equivalent like fractions and add them. You must use the LCD to get the answer correct. If possible, reduce the final sum. 1/5 2/3

OpenStudy (anonymous):

1/5 + 1/3 = 3/15 + 5/15 = 8/15

OpenStudy (anonymous):

thanks!! @kelliegirl33

OpenStudy (anonymous):

anytime :) glad to help